a school sold 200 tickets to a play. Ticket prices were $8.00 per adult and $5 per child. If the total sales for the tickets dcame to $1414, how many children's tickets were sold?
Answer by CharlesG2(834)

a school sold 200 tickets to a play. Ticket prices were $8.00 per adult and $5 per child. If the total sales for the tickets dcame to $1414, how many children's tickets were sold?


let A = adult tickets sold, and C = child tickets sold
A + C = 200 tickets
8A + 5C = $1414

C = 200 - A (now substitute this into 2nd equation)

8A + 5(200 - A) = 1414
8A + 1000 - 5A = 1414
8A - 5A = 414
3A = 414
A = 138 --> 138 adult tickets sold

138 + C = 200
C = 62 ----> 62 child tickets sold